onsider the following numerical example of the simple Keynesian model with no government spending, taxes or a foreign sector (all figures in R millions):
C = 100 + 0,9Y
I = 50
Answer the following questions.

A. Use a graph to illustrate the equilibrium level of output. [5]
1
Expert's answer
2019-08-15T14:18:16-0400

The equilibrium level is where the aggregate expenditure line intersects the 45 degree reference line. This occurs at Y1. 

"Y=AE=C+ I=100+0.9Y+50"

"0.1Y=150; Y=1500 R millions"

Thus Y1 is 1500R millions